OpenStudy (anonymous):

A batter hits a baseball upward with an initial speed of 96 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt - 16t2 where h represents height in feet and r represents the initial speed.

hero (hero):

h = rt - 16t^2 0 = 96t - 16t^2 0 = 16t(6 - t) t = 6

OpenStudy (saifoo.khan):

Everything is written in the question! D:

hero (hero):

@saifoo.khan , give me a medal

OpenStudy (saifoo.khan):

Give one to get one.

hero (hero):

Looks like you already have enough.

OpenStudy (saifoo.khan):

That dosnt matter..

hero (hero):

It matters.

OpenStudy (saifoo.khan):

No it dosnt.

OpenStudy (saifoo.khan):

8759 - 3975 = ?